# Motion Control via Metric-Aligning Motion Matching

### v1

@@ 3.2.5. Coarse-to-Fine Motion Alignment
Owing to the non-convex nature of GW problems (Solomon et al., 2016), proper initialization of the FSUGW block is crucial for guiding the optimization process toward high-quality aligned motion. To ensure effective initialization, we adopt a coarse-to-fine optimization strategy, as used in previous works (Li et al., 2022, 2023a; Elnekave and Weiss, 2022). This approach begins at a coarse level, with the initial transport plan $T$ computed by minimizing only the $L_{GW}$ term, and progressively refines the motion through upsampling until reaching the final stage:
@@
Input: control sequence $Y$, original motion $X$, parameters $\alpha$, $\lambda$, $\epsilon$, number of stages $N$, number of iterations $M$
@@
             $(\tilde{X^{\prime}},\tilde{Y},\tilde{X})\leftarrow$ ExtractPatches($X^{\prime},Y^{k},X^{k}$);
@@ 3.3.2. Other Constraints
Because a portion of our method relies on patch extraction and blending, it is compatible with various space-time control methods. In our implementation, we integrate hard keyframing and infinite loop constraints, as outlined in GenMM (Li et al., 2023a).
@@
When imposing an infinite loop constraint, the last patch of the aligned motion sequence is enforced to match the first patch of the same sequence, ensuring a seamless loop. Following the practice in GenMM (Li et al., 2023a), we achieve this by blending the overlapping region between the end and start patches consistently.
@@ 4. Applications
In this section, we demonstrate how the MAMM framework can be applied to various input control sequences. All use cases are realized by splitting each control sequence into patches, computing the distance matrix among the patches, and adjusting the optimization parameters; the underlying algorithm itself remains unchanged. All data used in the examples are sourced from Mixamo (Adobe, 2025) or Truebones (Truebones, 2025), which contain various motions and skeletons of humanoids and non-humanoids, respectively. The sequence lengths in our experiments range from approximately 50 to 600 frames at 30 fps. The patch size of the original motion was set to 11 frames, with a stride of one frame. Patch extraction from the control sequence used the same temporal duration and interval. Additionally, we used the AIST dance database (Tsuchida et al., 2019) for music data and the BEAT dataset (Liu et al., 2022) for speech data.
@@ 4.3. Motion-by-Numbers
Drawing inspiration from successful image synthesis pipelines guided by segmentation labels (Isola et al., 2017), we align motion to a user-specified temporal sequence of segmentation labels, each represented as a one-hot class vector. With MAMM, the original motion sequence does not require any precomputed or annotated segmentation labels; instead, the algorithm automatically aligns each segment according to the input labels. The user does not need prior knowledge of which label corresponds to which motion. However, when necessary, the user can explicitly specify the mapping by providing soft or hard constraints. Results are shown in the center row of Fig. 9 and Fig. 10, as well as in the supplemental video.
@@ 4.4. Motion Control by Audio
As mentioned in Sec. 2.3, audio-driven motion control has been extensively studied in the motion generation community. We focus on a scenario where the available example sequences are relatively short (3–20 s in duration) and lack paired audio or auxiliary annotations. The audio input is represented as 40-dimensional MFCC features, extracted using the Librosa library (McFee et al., 2024). Both speech and music data are used as audio controls, while short segments of gesture or dance motion serve as the original motion. As presented in Fig. 7, our method naturally synchronizes the resulting motion with style, beat, or volume changes in the audio, even when the original motion style differs from that in the audio. Notably, this approach does not require any training or specialized, hand-tuned optimization.
@@ 4.5. Motion-to-Motion Alignment
We applied MAMM to cross-skeletal motion-to-motion alignment, where one motion sequence is matched to another with a different skeletal structure (Li et al., 2024a). An example is shown in Fig. 8. Leveraging its metric-alignment principle, our method effectively handles both periodic motions, such as walking, and non-periodic activities, such as combat moves, without requiring additional training or customized feature engineering.

### v2

@@ 3.2.5. Coarse-to-Fine Motion Alignment
Owing to the non-convex nature of GW problems (solomon_entropic_2016), proper initialization of the FSUGW block is crucial for guiding the optimization process toward high-quality aligned motion. To ensure effective initialization, we adopt a coarse-to-fine optimization strategy, as used in previous works (li_ganimator_2022; li_example-based_2023; elnekave_generating_2022). This approach begins at a coarse level, with the initial transport plan $T$ computed by minimizing only the $L_{GW}$ term, and progressively refines the motion through upsampling until reaching the final stage:
@@
Input: control sequence $Y$, original motion $X$, parameters $\alpha$, $\lambda$, $\epsilon$, number of stages NN, number of iterations $M$
@@
[…]
@@ 3.3.2. Other Constraints
Because a portion of our method relies on patch extraction and blending, it is compatible with various space-time control methods. In our implementation, we integrate hard keyframing and infinite loop constraints, as outlined in GenMM (li_example-based_2023).
@@
When imposing an infinite loop constraint, the last patch of the aligned motion sequence is enforced to match the first patch of the same sequence, ensuring a seamless loop. Following the practice in GenMM (li_example-based_2023), we achieve this by blending the overlapping region between the end and start patches consistently.
@@ 4. Applications
In this section, we demonstrate how the MAMM framework can be applied to various input control sequences. All use cases are realized by splitting each control sequence into patches, computing the distance matrix among the patches, and adjusting the optimization parameters; the underlying algorithm itself remains unchanged. All data used in the examples are sourced from Mixamo (noauthor_mixamo_nodate) or Truebones (noauthor_truebones_nodate), which contain various motions and skeletons of humanoids and non-humanoids, respectively. The sequence lengths in our experiments range from approximately 50 to 600 frames at 30 fps. The patch size of the original motion was set to 11 frames, with a stride of one frame. Patch extraction from the control sequence used the same temporal duration and interval. Additionally, we used the AIST dance database (aist-dance-db) for music data and the BEAT dataset (liu2022beat) for speech data.
@@ 4.3. Motion-by-Numbers
Drawing inspiration from successful image synthesis pipelines guided by segmentation labels (isola_image--image_2017), we align motion to a user-specified temporal sequence of segmentation labels, each represented as a one-hot class vector. With MAMM, the original motion sequence does not require any precomputed or annotated segmentation labels; instead, the algorithm automatically aligns each segment according to the input labels. The user does not need prior knowledge of which label corresponds to which motion. However, when necessary, the user can explicitly specify the mapping by providing soft or hard constraints. Results are shown in the center row of Fig. 9 and Fig. 10, as well as in the supplemental video.
@@ 4.4. Motion Control by Audio
As mentioned in Sec. 2.3, audio-driven motion control has been extensively studied in the motion generation community. We focus on a scenario where the available example sequences are relatively short (3–20 s in duration) and lack paired audio or auxiliary annotations. The audio input is represented as 40-dimensional MFCC features, extracted using the Librosa library (mcfee_librosalibrosa_2024). Both speech and music data are used as audio controls, while short segments of gesture or dance motion serve as the original motion. As presented in Fig. 7, our method naturally synchronizes the resulting motion with style, beat, or volume changes in the audio, even when the original motion style differs from that in the audio. Notably, this approach does not require any training or specialized, hand-tuned optimization.
@@ 4.5. Motion-to-Motion Alignment
We applied MAMM to cross-skeletal motion-to-motion alignment, where one motion sequence is matched to another with a different skeletal structure (li_walkthedog_2024). An example is shown in Fig. 8. Leveraging its metric-alignment principle, our method effectively handles both periodic motions, such as walking, and non-periodic activities, such as combat moves, without requiring additional training or customized feature engineering.